OpenStudy (anonymous):

Calculate the slope of the line passing through the given points. If the slope is undefined, so state. Then indicate whether the line rises, falls, is horizontal, or is vertical. (8,2), (5,-3) (-3,4),(-3,5)

ganeshie8 (ganeshie8):

\[ slope = \frac{y2-y2}{x2-x1}\] slope is +ve : line rises slope is -ve : line falls slope is 0 : line is horizontal slope is 1/0 : line is vertical

ganeshie8 (ganeshie8):

(8,2), (5,-3) \[ slope = \frac{-3-2}{5-8} = \frac{-5}{-3} = \frac{5}{3}\] slope is +ve . so line rises